OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −1/2x − 2 y = 2x − 2 y = −1/2x − 7

OpenStudy (anonymous):

if tw lines r parallel then their slopes are equal so as slope of first line is 2 so general form y=mx+c now m=2 y=2x+c it passes through 5,3 put y=3 and x=5 3=10+c c=-7 so equation is y=2x-7 (a) :)

OpenStudy (anonymous):

Thanks a million mate =)